The length and width of a square are increased by 6 ft and 8 ft, respectively. The result is a rectangle whose area is 188 sq. ft more than the area of the square. Determine the length pf the side of the square.

Answer by ikleyn(52790)

.
(W+6)*(W+8) - W^2 = 188,

W^2 + 6W + 8W + 48 - W^2 = 188,

14W = 188-48,

14W = 140,

W = 10.

Answer. The side of the square is of 10 units long.